In a PARALLEL circuit, the potential difference across each resistor is *

1 point


inversely proportional to its resistance


directly proportional to its resistance


directly proportional to the square of its resistance


unrelated to its resistance


1
Expert's answer
2021-03-21T11:23:56-0400

In a parallel circuit, the voltage across each of the components is the same and the total current is the sum of the currents flowing through each component:


"V=V_1=V_2,""I=I_1+I_2."

Applying Ohm's Law, we get:


"I=\dfrac{V_1}{R_1}+\dfrac{V_2}{R_2}=\dfrac{V}{R_1}+\dfrac{V}{R_2},""I=V(\dfrac{1}{R_1}+\dfrac{1}{R_2})=\dfrac{V}{R_{eq}}."

Therefore, in a parallel circuit, the potential difference across each resistor is inversely proportional to its resistance.

Answer:

inversely proportional to its resistance